The ratio of the width of a rectangle to its length is 3:4. Write & solve a proportion to find the rectangle's width.
Answer by checkley77(12844)

W:L
3:4
W*L=108 IS THE AREA.
3N*4N=108
12N^2=108
N^2=108/12
N^2=9
N=SQRT9
N=3 IS THE MULTIPLIER.
3*3=9 WIDTH
3*4=12 LENGTH
PROOF:
9*12=108
108=108